OpenStudy (anonymous):

I really need help, I don't understand how to find marginal cost and revenue...

OpenStudy (amistre64):

marginal dhould bring derivatives to mind

OpenStudy (amistre64):

*should

OpenStudy (amistre64):

if cost = C(x), marginal cost is C'(x) if revenue = R(x), marginal revenue is R'(x)

OpenStudy (anonymous):

I don't exactly know what to do though. Is there a formula other than that.

OpenStudy (amistre64):

derivative rules ... unless you want to run thru first principles: given a function f(x), the marginal stuff can be calculated as\[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

So the function f(x) is what though? Like I f I have quantity, price, total revenue, and total cost. But I do not have marginal revenue, marginal cost, and profit? How do I fill everything out in the equation.

OpenStudy (amistre64):

f(x) is just a generic function notation for whatever you revenue function is (or cost function, or whatever function you are trying to marginalize)

OpenStudy (amistre64):

do you have a specific function that you would like to have marginalized?

OpenStudy (anonymous):

I have to make a graph. But for an example, I was giving quantity of 4, price was 38, total revenue was 152 and then total cost was 114.

OpenStudy (amistre64):

total cost should amount to: variable cost (times units sold) + fixed costs but the information you have provided is prolly taken out of some broader context.

OpenStudy (amistre64):

revenue - cost = gross income before taxes :/

OpenStudy (anonymous):

So can I still find what I am looking for with what I gave you?

OpenStudy (amistre64):

not really; there is no way i see with that scant information to develop any rule/expression to play with.

OpenStudy (amistre64):

if i we to ask you to find the slope of a line at the point (4,6); the solution is rather unobtainable simply because at a single point (and no expression for getting to that point) there is no way to develop the tolls necessary to calcualte it

OpenStudy (amistre64):

lets take the functions: y = x^2, y=x^3 ; the point (1,1) satisfies both equation; but the "marginal" results differ: y' = 2x, at x=1 is 2 y' = 3x^2 at x=1 is 3

OpenStudy (amistre64):

*to develop the tools ... not tolls :)

OpenStudy (anonymous):

I think I am getting it now.

OpenStudy (anonymous):

but what is the exponent 2 for?

OpenStudy (amistre64):

that was just some easy random functions to demonstrate that if all you are given is a point, there is no way to define the "marginal" aspects related to that point without a defininte expression

OpenStudy (amistre64):

different expressions yeild different results

OpenStudy (anonymous):

quantity- price total revenue marginal revenue total cost marginal cost profit 0 84 0 70 1 82 82 136 2 80 160 188 3 78 234 214 4 76 304 228 5 74 370 258 6 72 432 360 7 70 490 470 8 68 544 592

OpenStudy (anonymous):

that is what was given me

OpenStudy (anonymous):

the one all the way to the left is quantity, it gave me no profit

OpenStudy (amistre64):

the 70 column is total cost tho right?

OpenStudy (amistre64):

revenue = price, times quantity; you have a linear setup such that: price = 2q + 84 revenue = q(2q+84), or just 2q^2+84q marginal revenue would be based off of that setup

OpenStudy (amistre64):

price is 84-2q ... that is

OpenStudy (amistre64):

revenue = q(84-2q), or just 84q - 2q^2 you agree?

OpenStudy (anonymous):

Okayy, Yes I agree.

OpenStudy (amistre64):

then the marginal value for each q can be obtained from the expression: 84-4q if we were to run it thru that f(x+h)-f(x) limit definition i presented above

OpenStudy (amistre64):

the cost function is not wanting to be linear for us :/

OpenStudy (anonymous):

So what does that mean?

OpenStudy (amistre64):

that just means that trying to determine a suitable cost function is more difficult

OpenStudy (anonymous):

Great, just what I need. Lol I am so grateful you are helping me this much. You are a life saver!

OpenStudy (amistre64):

profit is just revenue minus cost ... so that column is pretty basic

OpenStudy (anonymous):

So profit doesn't have to do with marginal al all?

OpenStudy (amistre64):

marginal profit would :) but if its just labeled as profit ... then no; take the revenue and subtract the cost to get profit

OpenStudy (amistre64):

ima try to do an excel run for the cost .. make life simpler maybe

OpenStudy (amistre64):

this is what the cost data gives us as a graph, notice that it doesnt make a nice straight line

OpenStudy (amistre64):

well, the library decided to "upgrade" to excel 2013 ... which of course has to not have anything in the right places to do anything that you could do with excel 2010 with ....